OpenStudy (anonymous):

A large shipment of Granny Smith apples arrives at a grocery store. Weights of the apples are known to follow a normal distribution with mean 152 grams and standard deviation 16 grams.

OpenStudy (anonymous):

That is X~N(152,16^2) Then?

OpenStudy (anonymous):

(a) We take a random sample of three of the waiter's customers. Can you calculate the probability that the mean tip for these customers is less than $9.43? (b) If we take a random sample of 35 of the waiter's customers, what is the probability that the mean tip for these customers is greater than $6.13? (c) If we take a random sample of 45 of the waiter's customers, what is the probability that the mean tip for these customers is between $5.08 and $6.87?

OpenStudy (anonymous):

(a) np = mean If either or both np and nq are less than 5, a binomial distribution is to be used. Otherwise, a normal distribution is to be used. q=1-p p=mean/n Can you carry on from here?